# Balanced cross-Kerr coupling for superconducting qubit readout

### v1

@@ Abstract
We have shown that junction readout can realize a large cross-Kerr together with manageable resonator self-Kerr, as well as higher critical photon numbers compared to dispersive readout across a large resonator frequency range, all without the need for Purcell filters. We now present numerical simulations of junction readout suggesting that fast, high-fidelity, and quantum nondemolition readout of transmons can be achieved with realistic parameters.
Our simulations are based on integrating the stochastic Schrödinger equation for heterodyne measurements of the resonator first assuming an ideal readout efficiency $\eta=1$ [49], see [39] for further details. Because they use the full Hamiltonian Eq. 1, these simulations account for the cross-Kerr, $\chi_{z}/2\pi\simeq 10\>\textrm{MHz}$, but also the deleterious effect of the resonator self-Kerr, $K_{r}/2\pi\simeq 489\>\textrm{kHz}$. A two-step measurement pulse leading to an average photon number of $\bar{n}\simeq 40$ photons is used, well below the critical photon number of $n_{\rm{crit}}=65$. Despite this short integration time, we find a large separation of the measurement result distribution for the two initial state using an optimal discriminator (black dashed line), see Fig. 4(b). The resulting assignment error vs the integration time (full dark blue line) is reported in panel (c). For $\eta=1$, a measurement fidelity of 99.99% is obtained in as short as $t_{m}=20\>\textrm{ns}$. Furthermore, we find a QNDness of $99.89\%$ and $99.71\%$ for the ground and excited state, respectively [39].
[…]
@@
We have presented a circuit mediating a nonperturbative cross-Kerr coupling between a qubit and a readout resonator. The proposed junction readout method outperforms state-of-the-art dispersive readout by achieving higher critical photon numbers across a broad range of resonator frequencies and is less susceptible to the effect of gate charge. Notably, junction readout does not require Purcell filters and has comparable hardware overhead to dispersive readout. Even when accounting for finite qubit lifetimes and reduced readout efficiencies, this approach achieves a readout fidelity of $99.99\%$ in under $30$ ns. Junction readout overcomes many of the limitations posed by dispersive readout and can be implemented readily with only small modifications to current hardware. Junction readout could become a standard method for superconducting qubit measurement in next-generation quantum processors.
@@
During the preparation of this manuscript, we learned of similar work experimentally demonstrating Josephson junction-based readout of a transmon qubit. [51]. In contrast to our approach, in that work the resonator nonlinearity induced by the coupling junction is large such that a high-power bifurcation readout is used.
@@
The authors are grateful to Cristóbal Lledó for a critical reading of the manuscript. This work was undertaken thanks in part to funding from the U.S. Army Research Office Grant No. W911NF-23-1-0101. Additional support is acknowledged from NSERC, the Canada First Research Excellence Fund, and the Ministère de l’Économie et de l’Innovation du Québec.
@@
In this section, we outline the basic theory behind branch analysis, a numerical tool that can help predict the critical photon number for the onset of ionization. For a more detailed discussion of this approach, see Ref. [17, 21].
@@
Diagonalizing the above Hamiltonian results in a set of eigenvectors $\{|\lambda\rangle\}$. For each considered level $j_{t}$ of the transmon, we find the eigenvector from $\{|\lambda\rangle\}$ that maximizes the overlap $|\langle j_{t},0_{r}|\lambda\rangle|^{2}$ and label this eigenvector as $|\overline{j_{t},0_{r}}\rangle$, the overline denoting a dressed state. We then use these states $|\overline{j_{t},0_{r}}\rangle$ to recursively assign a label to all remaining states: |$\ket{\overline{j_{t},n_{r}+1}}$ is defined as the unassigned eigenstate $|\lambda\rangle$ which maximizes the overlap $|\langle\lambda|\hat{a}^{\dagger}|\overline{j_{t},n_{r}}\rangle|^{2}$. Repeating this for each transmon index $j_{t}$ results in a set of dressed states $B_{j_{t}}=\{\ket{\overline{j_{t},n_{r}}}\}$ which we refer to as branches.
Using these branches, we can compute the average transmon and resonator populations. An ionization process can be identified when the branches for the ground or excited state swaps with a branch of a higher excited state. The photon number at which this occurs is reported as the critical photon number.
S2 Extracting Purcell decay
In this section, we discuss how the Purcell-limited lifetimes $T_{1}^{\rm{Purcell}}$ shown in Fig. 1(c) are extracted. For each value of the charge-charge coupling $J$, we simulate a $T_{1}$ experiment by initializing the qubit in its dressed excited state |$\ket{\overline{1_{t},0_{r}}}$ and evolving it according to
@@
where $\hat{H}$ is given by Eq. 1. The transmon parameters are set to $E_{c}/2\pi=300\>\mathrm{MHz}$, $E_{J,\rm{total}}/E_{C}=50$, and the resonator parameters to $\omega_{r}/2\pi=9.375\>\mathrm{GHz}$, $Z_{r}=25\>\Omega$, and $\kappa/2\pi\simeq 2\chi_{z}/2\pi=8\>\mathrm{MHz}$. The coupling Josephson junction is characterized by $E_{J_{c}}/2\pi=4\>\mathrm{GHz}$.
In the above expression, the collapse operator $\hat{a}^{\prime}$ is expressed in the dressed basis as [52]
@@
where {|$\{\ket{\lambda}\}$ are the eigenstates of the transmon-resonator Hamiltonian $\hat{H}$ as defined above. Here, we consider only energy-loss processes, assuming the bath is in its vacuum state. To extract $T_{1}^{\mathrm{Purcell}}$, we fit the excited-state population |$\ket{\overline{1_{t},0_{r}}}\bra{\overline{1_{t},0_{r}}}$ | to a single exponential decay, $e^{-t/T_{1}^{\mathrm{Purcell}}}$.
The Purcell decay obtained from this method assumes a flat bath spectrum, i.e., $\kappa(\omega)=\kappa(\omega_{r})\equiv\kappa$. In practice, however, the qubit probes the bath at its own frequency, which for these parameters is $\overline{\omega}_{q}/2\pi=5.672$ GHz, rather than the resonator frequency. Therefore, to achieve a more accurate estimate, we normalize the $T_{1}^{\rm{Purcell}}$ value obtained above by a factor of $\kappa/\kappa(\overline{\omega}_{q})$  [53, 8, 54], which is approximately $(\overline{\omega}_{r}/\overline{\omega}_{q})^{2}\simeq 2.75$, assuming that all photon losses originate from coupling to the readout feedline [8, 54].
@@
Figure S1: Critical photon numbers for varying resonator frequencies $\omega_{r}$. The red dots indicate when there is no coupling capacitance, and the blue dots are for where the cancellation condition is met. Here the transmon parameters are the same as in Fig. 4 and the resonator impedance is $Z_{r}=35\>\Omega$. The coupling junction energy is $E_{Jc}/2\pi=7.8$ GHz, resulting in $\chi_{z}/2\pi\simeq 9$ MHz when the cancellation condition is met.
@@
As stated in the main text, the matrix element eliminated by the cancellation condition Eq. 3 is also precisely the matrix element responsible for transitioning the ground state of the transmon to the first excited state when measurement-induced transitions to higher excited states occur. Additionally, the matrix element that connects the first excited state to the second excited state is also approximately suppressed when the condition Eq. 3 is satisfied. Therefore, by eliminating these matrix elements the likelihood of multiphoton processes causing the transmon to jump in excitation is expected to be reduced.
@@
Fig. S1 shows the critical photon numbers for varying resonator frequencies, both for when the cancellation condition Eq. 3 is satisfied or not. When the cancellation condition is met, we observe as expected that the critical photon numbers increase.
@@
It is worth mentioning that for the dispersive readout, canceling the $0\leftrightarrow 1$ matrix element responsible for Purcell decay would also result in canceling most of the charge-charge coupling-based dispersive shift that is essential for readout [39]. Alternatively, as recently discussed in Ref. [55] in the context of dispersive readout, one could cancel parasitic matrix elements involving higher-energy states that are causing ionization without compromising the dispersive shift. However, no parameter set can cancel all parasitic matrix elements simultaneously. Furthermore, since these cancellations involve higher-energy states, they are sensitive to fluctuations in the gate charge. Consequently, this balanced coupling strategy does not push ionization to larger photon numbers for dispersive readout.
S3.2 Critical photon number threshold
In Fig. 3, we set the critical photon number threshold to be $n_{\rm{crit}}=35$. In this section, we show that the parameters chosen correspond to a physically meaningful readout performance metric. We begin with the analytical expression for the signal-to-noise ratio (SNR) in the long-time limit [56]
$\rm{SNR}\simeq\frac{2\epsilon}{\kappa}\sqrt{4\eta\kappa t_{m}}.$ (S4)
@@
In the above expression, $\epsilon$ is the drive amplitude, $\kappa$ is the resonator decay rate, $\eta$ is the measurement efficiency, and $t_{\mathrm{m}}$ is the measurement time. Assuming that the resonator is driven at its Lamb-shifted frequency, and for the optimal ratio $\kappa\simeq 2\chi$, which we ensure to be approximately satisfied in our numerical simulations, we can relate the above expression of $\rm{SNR}$ to the average photon number in the resonator $\bar{n}$ as
$\mathrm{SNR}\simeq 2\sqrt{2\eta\bar{n}\kappa t_{\mathrm{m}}},$ (S5)
@@
which simplifies to $\mathrm{SNR}\simeq 2\sqrt{2\eta\bar{n}}$ for a measurement time of order $t_{\mathrm{m}}\sim 1/\kappa$, corresponding to the resonator reaching its steady-state.
@@
Choosing the threshold to be $n_{\rm{crit}}=35$, we must ensure that the maximum number of photons placed in the resonator remains well below this threshold to minimize leakage. Given the approximately coherent-state nature of the resonator field during readout, as discussed in the main text, we model the photon number distribution as Poissonian and require its tail to have minimal overlap with the critical photon number. For instance, a safe metric is to limit this overlap to be approximately $0.01\%$. This condition translates to an average photon number $\bar{n}$ satisfying $n_{\rm{crit}}=\bar{n}+4\sqrt{\bar{n}}$, which yields $\bar{n}\simeq 18$.
@@
For a measurement efficiency of $\eta=0.5$, in line with state-of-the-art experiments [13, 20, 10], we find $\rm{SNR}\simeq 6$. The SNR can be linked to the measurement to the readout fidelity $F$ as [56]
@@
Note that this definition implies that $F\in[0.5,1]$. For $\rm{SNR}\simeq 6$, we find $F\simeq 0.9999(8)$, suggesting that the above choice of parameters and threshold comfortably lead to a readout fidelity exceeding $99.99\%$. Of course, the above expression does not account for distortions to the coherent states, transient dynamics of the readout, the $T_{1}$ limit of the transmon, and other effects. Nevertheless, they provide a useful and simple metric for relating readout performance to the critical photon number. In the main text and Sec. S5 we go beyond these simple approximations.
@@
The dispersive shift is often defined as $\chi_{z}=(E_{\overline{1,1}}-E_{\overline{1,0}}-E_{\overline{0,1}}+E_{\overline{0,0}})/2$, where $E_{\overline{i_{t},n_{r}}}$ denotes the energy of the dressed state |$\ket{\overline{i_{t},n_{r}}}$ of the transmon-resonator system, as labeled in Sec. S1. However, this definition of the dispersive shift is based on the low-energy spectrum of the transmon-resonator system and becomes inaccurate even for moderate photon numbers in the resonator [57, 58].
@@
for both dispersive readout and junction readout. At zero photons in the resonator, the parameters are taken such that the dispersive shift is $\chi_{z}/2\pi\simeq 10$ MHz for both readout methods. For dispersive readout, we find that at high photon numbers the magnitude of the dispersive shift is reduced by more than a half. Conversely, for junction readout the magnitude of the dispersive shift remains approximately constant. For example, at $n_{r}=35$, the dispersive shift is reduced to $\chi_{z}(n_{r}=35)/2\pi=-4.37$ MHz for dispersive readout, while $\chi_{z}(n_{r}=35)/2\pi=-8.526$ MHz for junction readout.
We attribute the significant difference between dispersive readout and junction readout to the nature of the coupling interaction. In the dispersive case, exact diagonalization of the Jaynes-Cummings Hamiltonian reveals that, beyond the linear dispersive regime, higher-order contributions to $\chi_{z}(n_{r})$ scales as $\sim(g/\Delta)^{2n_{r}}$ [59, 53, 60]. Conversely, for the $\cos\hat{\varphi}_{r}\cos\hat{\varphi}_{t}$ interaction, these higher-order contributions scale as $\varphi_{\text{zpfr}}^{2n_{r}}$ [38]. For the chosen set of parameters optimized to ensure high $n_{\rm{crit}}$ and $\chi_{z}$ for both readout methods, we find $\varphi_{\text{zpfr}}<|g/\Delta|$, and while $g/\Delta$ cannot be made arbitrarily small—otherwise, the leading-order dispersive shift becomes negligible— $\varphi_{\text{zpfr}}$ can be reduced by decreasing the resonator impedance, without sacrifying the leading-order dispersive shift.
@@
In addition, Fig. S2(b) shows the ratio between the matrix elements $|\langle\overline{0_{t},n_{r}+1}|\hat{H}_{\mathrm{JC,cancelled}}|\overline{1_{t},n_{r}}\rangle|$ and $|\langle\overline{0_{t},n_{r}+1}|\hat{H}_{\mathrm{JC}}|\overline{1_{t},n_{r}}\rangle|$ for varying resonator photon numbers. Here, $\hat{H}_{\mathrm{JC,cancelled}}=-E_{Jc}\sin{\hat{\varphi}_{t}}\sin{\hat{\varphi}_{r}}+J\hat{n}_{t}\hat{n}_{r}$ with $J$ chosen to satisfy Eq. 3, and $\hat{H}_{\mathrm{JC}}=-E_{Jc}\sin{\hat{\varphi}_{t}}\sin{\hat{\varphi}_{r}}$. We note that even at large resonator photon numbers, the ratio of the matrix elements remains small, showing that the cancellation condition remains effective during readout when the resonator is populated.
@@
In this section, we outline the details of our single-shot readout simulations. For the heterodyne readout simulations of Fig. 4, we use QuTiP’s solver [61, 62] to integrate the stochastic Schrödinger equation [49]
@@
Here, $\hat{H}$ is the sum of Eq. 1 and the drive term $\hat{H}_{d}=-i\epsilon(t)(\hat{a}-\hat{a}^{\dagger})$, where $\epsilon(t)$ is the resonator drive amplitude, with its explicit time dependence provided below. Additionally, $\kappa$ is the resonator decay rate, and $\hat{x}=\hat{a}+\hat{a}^{\dagger}$, and $\hat{p}=i(\hat{a}^{\dagger}-\hat{a})$ are the measured quadratures. In the above expression, the first and second lines describe the deterministic evolution of the state $\psi(t)$ while the last two terms describe the measurement back action. The two $dW_{i}$ are independent stochastic Wiener increments satisfying $\mathbb{E}[dW_{i}]=0$ and $\mathbb{E}[dW_{i}^{2}]=dt$. Both quadratures, $x$ and $p$, have their own independent Wiener increments, denoted by their respective subscripts.
For each simulated trajectory, we demodulate the signal with $e^{i\omega_{d}t}$ where $\omega_{d}$ is the drive frequency. The demodulated signal for the $x$ and $p$ quadratures are then integrated with the real and imaginary part of the optimal weight function $\langle a_{e}(t)\rangle-\langle a_{g}(t)\rangle$, respectively. The results are shown in Fig. 4(a) and (b). At each integration time we find the optimal discriminator in the IQ-plane which gives us the minimum assignment error defined as $\varepsilon=(P(e|g)+P(g|e))/2$, which is then shown in Fig. 4(c). We found that using nsubsteps = $12,000$, with the transmon dimension set to $7$ and resonator dimension set to $75$ showed good convergence.
Figure S3: (a) Resonator population as a function of integration time, for when the transmon is prepared in the ground or excited state. (b) and (c) show the population of the first $20$ energy levels of the transmon during the readout simulation, when starting in the ground and excited state, respectively. The circuit parameters are as given in Fig. 4.
As mentioned in the main text, we use a two-step pulse in order to fill the resonator with a large number of photons rapidly. The pulse shape is given as in Ref. [38],
@@
The results from the heterodyne readout simulations are compared to coherent state approximated assignment errors. Assuming the states in the resonator to be coherent states, we can get an expression for the optimal signal-to-noise ratio (SNR) as [44]
[…]
where $\eta$ is the measurement efficiency, $\kappa$ is the resonator decay rate, $t_{m}$ is the integration time, and $\alpha_{e/g}(t)$ is the resonator field. Using the above formula for the optimal SNR, we can compute the assignment error as [45, 12]

### v2

@@ Abstract
We have shown that junction readout can realize large cross-Kerr couplings together with small resonator self-Kerrs, higher critical photon numbers compared to dispersive readout across a wide range of resonator frequencies, and extended Purcell lifetimes. The ability to drive the resonator to large populations without encountering multiphoton resonances or significant leakage is crucial for achieving high-fidelity readout. In the following, we present numerical simulations of junction readout suggesting that fast, high-fidelity, and quantum nondemolition readout of transmons can be achieved with realistic parameters.
Our simulations are based on integrating the stochastic Schrödinger equation for heterodyne measurements of the resonator first assuming an ideal readout efficiency $\eta=1$ [Wiseman1993], see [Supplementary_Information] for further details. Because they use the full Hamiltonian Eq. 1, these simulations account for the cross-Kerr, $|\chi_{z}|/2\pi\simeq 10\>\textrm{MHz}$, but also the deleterious effect of the resonator self-Kerr, $|K_{r}|/2\pi\simeq 489\>\textrm{kHz}$. A two-step measurement pulse leading to an average photon number of $\bar{n}\simeq 40$ photons is used, well below the critical photon number of $n_{\rm{crit}}=65$. Despite this short integration time, we find a large separation of the measurement result distribution for the two initial state using an optimal linear discriminator (black dashed line), see Fig. 4(b). The resulting assignment error vs the integration time $t_{m}$ (full dark blue line) is reported in panel (c). For $\eta=1$, a measurement fidelity of 99.99% is obtained in as short as $t_{m}=20\>\textrm{ns}$. Furthermore, we find a QNDness of $99.89\%$ and $99.71\%$ for the ground and excited state, respectively [Supplementary_Information].
Because of the small resonator self-Kerr, the distortion of the qubit-state-dependent resonator coherent states is relatively small, see Fig. 4(a). Under this condition, the resonator state can be approximated as coherent, allowing us to compare our numerical results with the analytical expression for the signal-to-noise ratio (SNR) and the assignment error [Supplementary_Information, Gambetta:2007, RMP]. The assignment error derived from this approximation (full light blue line) closely matches the full stochastic simulations. The small discrepancy is primarily due to the self-Kerr nonlinearity of the resonator. Given the good agreement between the numerical simulation and the coherent-state approximation, we use the latter to estimate the readout fidelity for finite transmon lifetime and nonideal readout efficiency, something which would be otherwise numerically challenging. First, for $T_{1}=30\>\mu s$ (red triangles), the readout fidelity comfortably exceeds $99.9\%$ in $t_{m}\simeq 17\>\textrm{ns}$. For a longer-lived transmon with $T_{1}=120\>\mu s$ (green squares), the fidelity surpasses $99.99\%$ at $t_{m}\simeq 21\>\textrm{ns}$. Secondly, assuming an infinite $T_{1}$ but a readout efficiency of $\eta=0.5$ consistent with state-of-the-art experiments [Spring:2024, hazra_2024, Walter_disp_readout], we find a similar readout performance (dashed light blue line). Using quantum optimal control (QOC), the assignment error can be further reduced, enabling a readout fidelity of $99.99\%$ within $t_{m}\simeq 18$ ns using realistic pulse shapes. With QOC, the same fidelity can also be achieved in $75$ ns while including resonator reset [Supplementary_Information]. We emphasize that these results are based on parameters readily achievable with current hardware. Even accounting for reduced efficiency and $T_{1}$ limitations, our results indicate that junction readout could achieve a fidelity exceeding $99.99\%$ in under 30ns30\ $\text${ns} of integration time, outperforming state-of-the-art readout experiments by an order of magnitude.
@@
We have presented a circuit mediating a nonperturbative cross-Kerr coupling between a qubit and a readout resonator. The proposed junction readout method outperforms state-of-the-art dispersive readout by achieving higher critical photon numbers across a broad range of resonator frequencies and is less susceptible to the effect of gate charge. Notably, junction readout enhances the Purcell lifetime and has comparable hardware overhead to dispersive readout. Even when accounting for finite qubit lifetimes and reduced readout efficiency, this approach achieves a readout fidelity of $99.99\%$ in under $30$ ns of integration time. Junction readout overcomes many of the limitations posed by dispersive readout and can be implemented readily with only small modifications to current hardware. Junction readout could become a standard method for superconducting qubit measurement in next-generation quantum processors.
@@
During the preparation of this manuscript, we learned of similar work experimentally demonstrating Josephson junction-based readout of a transmon qubit [wang:2024]. In contrast to our approach, in that work the resonator nonlinearity induced by the coupling junction is large such that a high-power bifurcation readout is used.
@@
The authors are grateful to Cristóbal Lledó for a critical reading of the manuscript, and to Peter Spring and Geneviève Gervais for insightful discussions. This work was undertaken thanks in part to funding from the U.S. Army Research Office Grant No. W911NF-23-1-0101. Additional support is acknowledged from NSERC, the Canada First Research Excellence Fund, and the Ministère de l’Économie et de l’Innovation du Québec.
@@
We outline the basic theory behind branch analysis, a numerical tool used to predict the critical photon number for the onset of ionization. For a more detailed discussion of this approach, see Refs. [Boissonneault_qubit_readout, Dynamics_of_transmon_ionzation, Dumas2024].
@@
[…]
Using these branches, we can compute the average transmon and resonator populations. An ionization process can be identified when the branches for the ground or excited state swaps with a branch of a higher excited state. We define the critical photon number for the ground (excited) state to be when the transmon population exceeds $N_{t}=2\>(3)$ for at least 3 photons in a row, see Refs. [Dumas2024, chapple:2024] for further details. The definition of the ionization critical photon number used here was shown to agree well with experimental results [Dumas2024].
S2 Extracting Jaynes–Cummings induced Purcell decay
[…]
@@
where $\hat{H}$ is given by Eq. (1). The transmon parameters are set to $E_{C}/2\pi=300\>\mathrm{MHz}$, $E_{J,\rm{total}}/E_{C}=50$, and the resonator parameters to $\omega_{r}/2\pi=9.375\>\mathrm{GHz}$, $Z_{r}=25\>\Omega$, and $\kappa/2\pi\simeq 2|\chi_{z}|/2\pi=8\>\mathrm{MHz}$. The coupling Josephson junction is given by $E_{J_{c}}/2\pi=4\>\mathrm{GHz}$.
In the above expression, the collapse operator $\hat{a}^{\prime}$ is expressed in the dressed basis as [BRE:2002]
@@
[…]
The Purcell decay obtained from this method assumes a flat bath spectrum, i.e., $\kappa(\omega)=\kappa(\omega_{r})\equiv\kappa$. In practice, however, the qubit probes the bath at its own frequency, which for these parameters is $\overline{\omega}_{q}/2\pi=5.672$ GHz, rather than the resonator frequency. Therefore, to achieve a more accurate estimate, we normalize the $T_{1}^{\rm{Purcell}}$ value obtained above by a factor of $\kappa/\kappa(\overline{\omega}_{q})$  [Boissonneault_dressed_dephasing, RMP, Labarca2024], which is approximately $(\overline{\omega}_{r}/\overline{\omega}_{q})^{2}\simeq 2.75$, assuming that all photon losses originate from coupling to the readout feedline [RMP, Labarca2024]. Finally, since the Jaynes–Cummings–induced Purcell decay $T_{1}^{\text{Purcell}}$ scales as $\sim\kappa\times(g/\Delta)^{2}$, with $g$ the effective Jaynes–Cummings coupling that itself scales linearly with $J$ as well as with the circuit parameters $E_{J_{c}}$ and $\sqrt{\omega_{q}}$ [RMP], varying these parameters leads to the same conclusion.
@@
Figure S1: Critical photon numbers for varying resonator frequencies $\omega_{r}$. The red dots indicate when there is no coupling capacitance, and the blue dots are for where the cancellation condition is met. Here the transmon parameters are the same as in Fig. 4 and the resonator impedance is $Z_{r}=35\>\Omega$. The coupling junction energy is $E_{Jc}/2\pi=7.8$ GHz, resulting in $|\chi_{z}|/2\pi\simeq 9$ MHz when the cancellation condition is met.
@@
As stated in the main text, the matrix element eliminated by the cancellation condition Eq. (3) is also precisely the matrix element responsible for transitioning the ground state of the transmon to the first excited state when measurement-induced transitions to higher excited states occur. Additionally, the matrix element that connects the first excited state to the second excited state is also approximately suppressed when the condition Eq. (3) is satisfied. Therefore, by eliminating these matrix elements the likelihood of multiphoton processes causing the transmon to jump in excitation is expected to be reduced.
@@
Fig. S1 shows the critical photon numbers for varying resonator frequencies, comparing when the cancellation condition Eq. (3) is satisfied with the no-cancellation case ($J=0$). When the cancellation condition is met, we observe as expected that the critical photon numbers increase. We also find that canceling the $1\leftrightarrow 2$ matrix element, rather than the $0\leftrightarrow 1$ element as specified in Eq. (3), yields essentially the same result: the optimal coupling $J$ and the corresponding critical photon numbers remain essentially unchanged.
@@
It is worth mentioning that for the dispersive readout, canceling the $0\leftrightarrow 1$ matrix element would also result in canceling most of the charge-charge coupling-based dispersive shift that is essential for readout. Alternatively, as recently discussed in Ref. [Sank:2024] in the context of dispersive readout, one could cancel parasitic matrix elements involving higher-energy states that are causing ionization without compromising the dispersive shift. However, no parameter set can cancel all parasitic matrix elements simultaneously. Furthermore, since these cancellations involve higher-energy states, they are sensitive to fluctuations in the gate charge. Consequently, this balanced coupling strategy does not push ionization to larger photon numbers for dispersive readout.
S3.2 Critical photon number percentile
Figure S2: Critical photon number at the 5th percentile across resonator frequencies, defined as the value of $n_{\rm crit}$ that $95\%$ of realizations exceed, for both junction readout (blue) and dispersive readout (red). Transmon parameters are the same as in Fig. (3). Dashed horizontal line indicates $n_{\rm{crit}}=30$.
We further compare dispersive and junction readout in terms of their critical photon numbers, and discuss how a larger critical photon number enables higher-fidelity readout. Fig. S2 shows the $5$th percentile critical photon number for each resonator frequency, defined as the value of $n_{\rm crit}$ that $95\%$ of realizations exceed, for both junction readout (blue) and dispersive readout (red). For all data points and both readout schemes, the dispersive shift is fixed at $|\chi_{z}|/2\pi\simeq 9$ MHz. As noted in the main text, junction readout exhibits a broad range of resonator frequencies where $95\%$ of the critical photon numbers exceed $30$ for random transmon gate charges, whereas no such range exists for dispersive readout. With this threshold of $n_{\rm crit}=30$ and the parameters $\chi_{z}$ and $K_{r}$ reported in the main text, high performance readout can be achieved.
To see that, we begin with the analytical expression for the signal-to-noise ratio (SNR) in the long-time limit [Gambetta2008],
$\rm{SNR}\simeq\frac{2\epsilon}{\kappa}\sqrt{2\eta\kappa t_{m}}.$ (S4)
@@
In the above expression, $\epsilon$ is the drive amplitude, $\kappa$ is the resonator decay rate, $\eta$ is the measurement efficiency, and $t_{\mathrm{m}}$ is the measurement time. Assuming that the resonator is driven at its Lamb-shifted frequency, and for the optimal ratio $\kappa\simeq 2|\chi_{z}|$, which we ensure to be approximately satisfied in our numerical simulations, we can relate the above expression of $\rm{SNR}$ to the average photon number in the resonator $\bar{n}$ as
$\mathrm{SNR}\simeq\sqrt{4\eta\bar{n}\kappa t_{\mathrm{m}}},$ (S5)
@@
which simplifies to $\mathrm{SNR}\simeq 2\sqrt{2\eta\bar{n}}$ for a measurement time of order $t_{\mathrm{m}}\sim 1/|\chi_{z}|$, corresponding to the resonator reaching its steady-state [Walter_disp_readout].
@@
Choosing the threshold to be $n_{\rm{crit}}=30$, we must ensure that the maximum number of photons placed in the resonator remains well below this threshold to minimize leakage. Given the approximately coherent-state nature of the resonator field during readout, as discussed in the main text, we model the photon number distribution as Poissonian and require its tail to have minimal overlap with the critical photon number. For instance, a safe metric is to limit this overlap to be approximately $0.01\%$. This condition translates to an average photon number $\bar{n}$ satisfying $n_{\rm{crit}}=\bar{n}+4\sqrt{\bar{n}}$, which yields $\bar{n}\simeq 14.7$.
@@
For a measurement efficiency of $\eta=0.5$, in line with state-of-the-art experiments [Spring:2024, hazra_2024, Walter_disp_readout], we find $\rm{SNR}\simeq 7.66$. The SNR can be linked to the measurement to the readout fidelity $F$ as [Gambetta2008]
@@
Note that this definition implies that $F\in[0.5,1]$. For $\rm{SNR}\simeq 7.66$, we find $F\simeq 0.9999(4)$, suggesting that the above choice of parameters and threshold comfortably lead to a readout fidelity exceeding $99.99\%$. Of course, the above expression does not account for distortions to the coherent states, transient dynamics of the readout, the $T_{1}$ limit of the transmon, and other effects. Nevertheless, they provide a useful and simple metric for relating readout performance to the critical photon number. In the main text and Sec. S5 we go beyond these simple approximations.
@@
[…]
@@
for both dispersive readout and junction readout. At zero photons in the resonator, the parameters are taken such that the dispersive shift is $|\chi_{z}|/2\pi\simeq 10$ MHz for both readout methods. For dispersive readout, we find that at high photon numbers the magnitude of the dispersive shift is reduced by more than a half. Conversely, for junction readout the magnitude of the dispersive shift remains approximately constant. For example, at $n_{r}=35$, the dispersive shift is reduced to $\chi_{z}(n_{r}=35)/2\pi=-4.37$ MHz for dispersive readout, while $\chi_{z}(n_{r}=35)/2\pi=-8.526$ MHz for junction readout.
We attribute the significant difference between dispersive readout and junction readout to the nature of the coupling interaction. In the dispersive case, exact diagonalization of the Jaynes-Cummings Hamiltonian reveals that, beyond the linear dispersive regime, higher-order contributions to $\chi_{z}(n_{r})$ scales as $\sim(g/\Delta)^{2n_{r}}$ [Boissonneault_qubit_readout, Boissonneault_dressed_dephasing, Boissonneault_qubit_dephasing]. Conversely, for the $\cos\hat{\varphi}_{r}\cos\hat{\varphi}_{t}$ interaction, these higher-order contributions scale as $\varphi_{\text{zpfr}}^{2n_{r}}$ [chapple:2024]. For the chosen set of parameters optimized to ensure high $n_{\rm{crit}}$ and $\chi_{z}$ for both readout methods, we find $\varphi_{\text{zpfr}}<|g/\Delta|$, and while $g/\Delta$ cannot be made arbitrarily small—otherwise, the leading-order dispersive shift becomes negligible— $\varphi_{\text{zpfr}}$ can be reduced by decreasing the resonator impedance, without sacrificing the leading-order dispersive shift.
@@
In addition, Fig. S3(b) shows the ratio between the matrix elements $|\langle\overline{0_{t},n_{r}+1}|\hat{H}_{\mathrm{JC,cancelled}}|\overline{1_{t},n_{r}}\rangle|$ and $|\langle\overline{0_{t},n_{r}+1}|\hat{H}_{\mathrm{JC}}|\overline{1_{t},n_{r}}\rangle|$ for varying resonator photon numbers. Here, $\hat{H}_{\mathrm{JC,cancelled}}=-E_{Jc}\sin{\hat{\varphi}_{t}}\sin{\hat{\varphi}_{r}}+J\hat{n}_{t}\hat{n}_{r}$ with $J$ chosen to satisfy Eq. (3), and $\hat{H}_{\mathrm{JC}}=-E_{Jc}\sin{\hat{\varphi}_{t}}\sin{\hat{\varphi}_{r}}$. We note that even at large resonator photon numbers, the ratio of the matrix elements remains small, showing that the cancellation condition remains effective during readout when the resonator is populated.
@@
In this section, we outline the details of our single-shot readout simulations. For the heterodyne readout simulations of Fig. 4, we use QuTiP’s solver [Qutip_1, Qutip_2] to integrate the stochastic Schrödinger equation [Wiseman1993]
@@
[…]
For each simulated trajectory, we demodulate the signal with $e^{i\omega_{d}t}$. The demodulated signal for the $x$ and $p$ quadratures are then integrated with the real and imaginary part of the optimal weight function $\langle a_{e}(t)\rangle-\langle a_{g}(t)\rangle$, respectively. The results are shown in Fig. 4(a) and (b). At each integration time we find the optimal discriminator in the IQ-plane which gives us the minimum assignment error defined as $\varepsilon=(P(e|g)+P(g|e))/2$, which is then shown in Fig. 4(c). We found that using nsubsteps = $12,000$, with the transmon dimension set to $7$ and resonator dimension set to $75$ showed good convergence. We confirm with branch analysis that this truncation of the Hilbert space does not alter the ionization critical photon number or the dispersive shift.
Figure S4: (a) Resonator population as a function of integration time, for when the transmon is prepared in the ground or excited state. (b) and (c) show the population of the first 2020 energy levels of the transmon during the readout simulation, when starting in the ground and excited state, respectively. The circuit parameters are as given in the caption of Fig. 4.
As mentioned in the main text, we use a two-step pulse in order to fill the resonator with a large number of photons rapidly. The pulse shape is given as in Ref. [chapple:2024],
@@
The results from the heterodyne readout simulations are compared to coherent state approximated assignment errors. Assuming the states in the resonator to be coherent states, we can get an expression for the optimal signal-to-noise ratio (SNR) as [Bultink2018]
$\mathrm{SNR}(t_{\mathrm{m}})=\sqrt{2\kappa\eta\int_{0}^{t_{\mathrm{m}}}|\alpha_{e}(t)-\alpha_{g}(t)|^{2}\>dt},$ (S10)
where $\eta$ is the measurement efficiency, $\kappa$ is the resonator decay rate, $t_{m}$ is the integration time, and $\alpha_{e/g}(t)$ is the resonator field. Using the above formula for the optimal SNR, the assignment error is bounded by [Gembetta_optimal_readout, swiadek2023enhancing]